Schopp Corporation makes a mechanical stuffed alligator that sings the Martian national anthem. The following information is available for Schopp Corporation's anticipated annual volume of 508,000 units.

Per Unit Total

Direct materials $ 7

Direct labor $11

Variable manufacturing overhead $17

Fixed manufacturing overhead $3,048,000

Variable selling and administrative expenses $16

Fixed selling and administrative expenses $1,524,000

The company has a desired ROI of 25%. It has invested assets of $30,480,000.

1.Compute the total cost per unit.

2. Compute the desired ROI per unit.

3. Compute the markup percentage using total cost per unit. (%)

4. Compute the target selling price.
Answer #1

Total cost per unit

= 7+11+17+16+[(3,048,000+1,524,000)/508,000]

= 60

Desired ROI per unit = (30,480,000*25%)/508,000

= 15

Markup percentage = 15/60

= 25%

Tarhet selling price = 60 + 15

= 75